Suppose you roll a special 6-sided die. What is the probability that the number rolled is a "4" OR a "5"? Enter you answer as a fraction or a decimal with 3 decimal digits.
Answer #1

here total number of outcome on a  6-sided die =6

number of favorable outcome =2 (4 and 5)

therefore  probability that the number rolled is a "4" OR a "5" =favorable outcome/total outcome

=2/6 =1/3